OpenStudy (anonymous):

how do you find the slope of the tangent line for: y^2=x^3/(4-x)

OpenStudy (anonymous):

at the point (2,2)

OpenStudy (ranga):

Do implicit differentiation and evaluate y' at (2,2)

OpenStudy (anonymous):

Do you take the square root first?

OpenStudy (ranga):

No need to take square root. y^2=x^3/(4-x) differentiate both sides: 2yy' = { (4-x)(3x^2) _ x^3(-1) } / (4-x)^2 put x = 2, y = 2 above and solve for y'

OpenStudy (anonymous):

Ohhhhh right, okay! Thank you!

OpenStudy (ranga):

you are welcome.